Assume that you are considering purchasing stock as an investment. You have narrowed the choice to either Superiority Corporation stock or Internet Company stock and have assembled the following data for the two companies. Your strategy is to invest in companies that have low​ price-earnings ratios but appear to be in good shape financially. Assume that you have analyzed all other factors and that your decision depends on the results of ratio analysis.1. Compute the following ratios for both companies for the current year, and decide which companys stock better fits your inSelected income statement data for the current year: Internet Superiority 599,000 $ 516,000 Net sales (all on credit). Cost oSelected balance sheet and market price data at end of current year: Superiority Internet Current assets: Cash 24,000 $ 37,00Selected balance sheet data at beginning of current year: Superiority Internet Balance sheet: Current receivables, net 144,00
Answer #1
Superiority Internet
a Quick ratio =(current assets-inventories-prepaid expenses)/current liabilities 0.62 0.63
=(451000-210000-18000)/361000 =(414000-185000-13000)/341000
b Inventory trunover =cost of goods sold/average invenotries                                                            2.19                                                            1.99
=456000/((210000+206000)/2) =382000/((185000+198000)/2)
c Days sales in average receivables =Average accounts receivable*365/total credit sales                                                        101.46                                                        126.62
=(((189000+144000)/2)*365)/599000 =(((162000+196000)/2)*365)/516000
d Debt ratio Total liabilities/total assets                                                            0.69                                                            0.73
=670000/977000 =687000/935000
e Times interest earned ration EBIT/interest expenses - 4.75
=76000/16000
f Return on common equity =(net income-preferred dividend)/average common equity                                                            0.53                                                            0.78
=61000/115000 =(40000-1200)/50000
g EPS =(net income-preferred dividend)/no. of shares equity                                                            0.53 3.88
=(61000)/115000 =(40000-1200)/10000
h PE ratio price/eps 10                                                            4.97
=5.3/0.53 =3.88/0.78
Internet company's PE ratio has low PE ration and is financially fit accoriding to all ration. So Interment company should be selected.